OpenStudy (anonymous):

Identify the y-intercept. y = - 5/8 x + 5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

y=5/8(0)+5 =5

OpenStudy (aroub):

how? because, make x zero y = - 5/8 x + 5 y=-5/8 (0)+5 y=5

OpenStudy (anonymous):

why is it zero ?

OpenStudy (aroub):

or.. because its in the form y=mx+b b is always the y-intercept so in here its 5.